The demand for ice cream cones is P=1600 and Qd is 2
The supply of ice cream cones is P =400 and Qs is 1. The price of a cone is expressed in cents, and the quantities are expressed in cones per day. To find the equilibrium price (P*) and the equilibrium quantity (Q*), substitute Q* for QD and QS and P* for P. (8 marks). (FOR QD, P=a-bQd and for QS, P=c+dQS).
1
Expert's answer
2020-04-20T18:46:15-0400

Equilibrium point is the point at which the market will clear. The demand function based on the above question will be:

P=1600-2Q : The demand function has a negative slope.

The supply function will be as follows:

P=400+Q : The supply function has a positive slope.

The equilibrium point is the point where the quantity demanded and the quantity supplied are equal.

Therefore, equilibrium condition is as follows:

1600-2Q=400+Q

1600-400=Q+2Q

1200=3Q

Q=1200/3

Q*=400

Equilibrium price will therefore be,

P=400+Q

P=400+400

P*=800

OR

P=1600-2Q

P=1600-2*400

P*=800

Substituting with Q* for Qd and P with P* get:

P=a-2bQd

800=a-800 as the demand equation.

a=1600


P=c+dQS

800=c+400 as the supply function.

c=400